OpenStudy (anonymous):

What is the difference of (8.5 × 10^15) – (5.0 × 10^14) in correct scientific notation? A. 3.5 × 10^1 B. –5.15 × 10^14 C. 8.0 × 10^15 D. 3.5 × 10^15 E. 9.0 × 10^14

OpenStudy (anonymous):

Is A the correct answer?

OpenStudy (anonymous):

I subtracted 8.5 - 5.0 and got 3.5

OpenStudy (anonymous):

and then 10^15 - 10^14 = 10^1

thomaster (thomaster):

you can write this as 8.5 - 0.5 because the exponent of the 5.0 is 1 lower than the 8.5 when you do \(1.5*10^3 - 5.0*10^2\), it's the same as 1500 = 500. So the answer will be \(1.0*10^3\) Do you have an idea what might be the answer?

OpenStudy (anonymous):

8.0 x 10^ 15?

thomaster (thomaster):

yes :)

OpenStudy (anonymous):

Thank you so much!